What is the wavelength in air under standard conditions of a compressional wave

whose frequency is 250 per second? Assume that the bulk modulus for air is 1.40

x 106 dyn/cm2 and the density of air is 1.29 g/L.


1
Expert's answer
2022-03-02T10:49:21-0500

The speed of sound

"v=\sqrt{\frac{B}{\rho}}=\sqrt{\frac{1.4*10^5}{1.29}}=329\:\rm m/s"

The wavelength

"\lambda=\frac{v}{f}=\frac{329}{250}=1.32\:\rm m"